A single fair 6-sided die is rolled, and the events E and F are as follows: E={An even number shows} F={A number greater than 3 shows}. Find “the probability of event E OR event F occurring”, P(E U F).
Answer by bhoopal624(2)

events of E={2,4,6} =>p(E)=3/6=1/2
events of F={4,5,6} =>p(F)=3/6=1/2
common events of E and F ={4,6} =>p(c)=2/6=1/3
P(E U F)=(1/2)+(1/2)-(1/3)
=2/3